EVA Report – December 30th

Crew 236 EVA Report 30-DEC-2021

EVA # 11

Author: Vladimir Zeltsman

Purpose of EVA: Evaluation of soil and ground conditions for additional potential habitat locations and a geology survey of the Summerville formation/layers

Start time: 9:00 am

End time: 13:30 pm

Narrative: The EVA was originally requested for near and east of the summer view overlook, but changed to Candor Chasma. The coordinates are originally N4253000 W522800 but rescheduled for N4251000 W521000. The crew entered the airlock at 9:15 and headed towards Candor Chasma shortly after. Once both rovers were parked on Galileo, the crew headed east toward the destination. After hiking down into Candor Chasma, 1 location was measured to see if it was large enough for a habitat. The one location was open and had acceptable ground and soil conditions but significant terraforming would be required to make a road and clear the land of all fauna. This leads Vladimir (GHO) to believe that further east into Candor Chasma, there could be locations large and solid enough for a potential habitat location. At 11:15, the crew turned around and headed back to the rovers due to two suit failures. The crew arrived back at the Hab at 12:26. The EVA, although not producing the desired results, is still considered a success, since multiple sites were evaluated.

Destination: Candor Chasma/Region G

Coordinates (use UTM NAD27 CONUS): N4251000, W521000

Participants: Cesare (XO), Vladimir (GHO), Tyler (SCI), and Ben (JOU)

Road(s) and routes per MDRS Map: From the Hab, driving north on Cow Dung Rd to Galileo Rd and Galileo Rd to near area F; walking from area F east to Candor Chasma

Mode of travel: Driving and walking

EVA Report – December 28th

Crew 236 EVA Report 28-DEC-2021

EVA # 9

Author: Kasey Hilton

Purpose of EVA: Evaluation of soil and ground conditions for additional potential habitat locations

Start time: 10:30

End time: 14:00

Narrative: The crew entered the airlock at 10:35 and headed towards Candor Chasma shortly after. Once both rovers were parked on Galileo, the crew headed east toward the destination. After hiking down into Candor Chasma, 2 locations were measured to see if they were large enough for a habitat, but neither were, however the ground and soil conditions were acceptable. This leads Vladimir (GHO) to believe that further east into Candor Chasma, there could be locations large and solid enough for a potential habitat location. At 12:15, the crew turned around and headed back to the rovers. The crew arrived back at the Hab at 12:54. The EVA, although not producing the desired results, is still considered a success, since multiple sites were evaluated.

Destination: Candor Chasma/Region G

Coordinates (use UTM NAD27 CONUS): N4251000, W521000

Participants: Kasey (CMD), Vladimir (GHO), and Ben (JOU)

Road(s) and routes per MDRS Map: From the Hab, driving north on Cow Dung Rd to Galileo Rd and Galileo Rd to near area F; walking from area F east to Candor Chasma

Mode of travel: Driving and walking

EVA Report – December 20th

EVA #1

Author: Cesare Guariniello

Start time: 11.38 AM

End time: 12.37 PM

Narrative: Team conducted a training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. The crew explored the hills around Marble ritual and collected rock samples before heading back to the hab.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: XO, ASTRO, GHO, SCI

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Curiosity

EVA #2

Author: Cesare Guariniello

Start time: 2.04 PM

End time: 3.18 PM

Narrative: Team conducted a training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. The crew explored the hills around Marble ritual and collected rock samples before heading back to the hab.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: CMD, ENG, JOU

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Opportunity

Cesare Guariniello

XO/HSO

Crew 236

Sol Summary – December 28th

Crew 236 Sol Summary Report 28-DEC-2021

Sol: 9

Summary Title: One Hab to Rule Them All

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: The crew awoke at a modest 8:00 and by 9:00 was eating crepes with Nutella and sweetened condensed milk. Not a bad way to start the morning. By 10:00, it was time to start preparing for today’s EVA, which took Vladimir (GHO), Ben (JOU), and I to Candor Chasma in search of alternate habitat locations. Our descent into Candor Chasma felt a lot like Bilbo Baggins’s trek into the goblin caves in the Misty Mountains. Except it wasn’t dark. And there weren’t any goblins. And instead of a ring, there was a potential location for another Hab hidden somewhere. But besides that, it was the same. There was even a weird, pale creature following us the whole way! I guess you could call him Gollum, but we normally just call him Ben… Anyways, after EVA, the crew fell into their same routine of a short nap and continuing to chip away at research and outreach projects. After a dinner, we plan to watch the rest of Apollo 13!

Look Ahead Plan: The first half of tomorrow will be spent starting to wrap up research and outreach projects and then there’s an EVA scheduled for the afternoon. The crew on EVA will be going back to Barianca Butte for additional thermal readings and scouting for additional Hab locations.

Anomalies in work: None

Weather: Overcast, windy, with temperatures in the 30s

Crew Physical Status: Health and Safety Officer, Cesare, yelled “good!” from downstairs, so we must be in pretty good shape!

EVA: EVA to Candor Chasma was successful although it did not yield the results for Vladimir’s (GHO) project on finding alternate habitat locations as hoped.

Reports to be filed: Operations Report, GreenHab Report, EVA Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

Sol Summary – December 30th

Crew 236 Sol Summary Report 30-DEC-2021

Sol: 11

Summary Title: The One Where They Leave Mars

Author’s name: Kasey Hilton

Mission Status: Nominal, last day on the fourth rock from the Sun

Sol Activity Summary: Today was the last day of sim. It’s bittersweet. On one hand, we dreamed of hot tubs and gin and tonics over our oatmeal breakfast, while on the other, we reminisced over the past EVAs while waiting to depressurize in the airlock. Once the crew left for the last EVA, Dylan (ASTRO) spent some time in the Musk Observatory, while Pavi (ENG) and I chatted about life sitting by the downstairs heater. After EVA and chowing down on some hot soup, the crew gathered around the kitchen table and took turns painting each other’s nails a cheddar cheese color, galaxy, or a combination of the two. The remainder of the afternoon was spent finishing up some research and outreach. Cesare (XO) baked some rocks to analyze the water content in the ScienceDome, I filmed a video answering questions from a third grade class, and some others studied the back of their eyelids. Our last dinner in the Hab was homemade pizza made by our resident Italian and we’ll probably spend the next few hours in a heated debate about which movie we should watch as our last on Mars. When we wake up, we’ll be back on Earth!

Look Ahead Plan: We will be out of sim tomorrow at 8am. Majority of tomorrow will be spent cleaning the Hab, completing the food inventory, and shedding a few tears when thinking of leaving the place we’ve called Hab, sweet Hab, for the last two weeks.

Anomalies in work: None

Weather: Partly cloudy, temperatures in the low 40s F

Crew Physical Status: Nominal

EVA: Today’s EVA was originally planned for Summerville Overlook, but was rescheduled to Candor Chasma. The EVA ended early due to two space suit fans dying, but was still considered successful because Cesare (XO) finding four rocks composed of salts and Vladimir (GHO) found one potential Hab location.

Reports to be filed: Operations Report, GreenHab Report, EVA Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

Sol Summary – December 21st

Crew 236 Sol Summary Report 21-DEC-2021

Sol: 2

Summary Title: A Slow Day

Author’s name: Kasey Hilton

Mission Status: Nominal

Sol Activity Summary: With no EVA scheduled for the day, the crew had a late wake up finished off with pancakes and Nutella for breakfast. With full bellies, everyone spent the afternoon working on research and outreach projects, processing images captured in the observatories, and harvesting a long awaited carrot and cucumber. By the end, everyone felt refreshed and recharged. Today was indeed a slow day in the Hab for crew 236.

Look Ahead Plan: There is an EVA planned for tomorrow for collecting rock samples. Everyone else at the Hab will be continuing to work on research and outreach from inside the Hab.

Anomalies in work: None

Weather: Partly overcast, temperature in the 30s

Crew Physical Status: Everyone is happy and healthy. However, crew journalist, Ben, is quoted saying that he “cannot be defeated by conventional weapons”, but I guess a metal stair to the shin does not fall into that classification. No physical injury occurred, just an injury to his pride and a short moment of pain.

EVA: No EVA today

Reports to be filed: Operations Report, GreenHab Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

Operations – December 28th

Crew 236 EVA Report 28-DEC-2021

EVA # 9

Author: Kasey Hilton

Purpose of EVA: Evaluation of soil and ground conditions for additional potential habitat locations

Start time: 10:30

End time: 14:00

Narrative: The crew entered the airlock at 10:35 and headed towards Candor Chasma shortly after. Once both rovers were parked on Galileo, the crew headed east toward the destination. After hiking down into Candor Chasma, 2 locations were measured to see if they were large enough for a habitat, but neither were, however the ground and soil conditions were acceptable. This leads Vladimir (GHO) to believe that further east into Candor Chasma, there could be locations large and solid enough for a potential habitat location. At 12:15, the crew turned around and headed back to the rovers. The crew arrived back at the Hab at 12:54. The EVA, although not producing the desired results, is still considered a success, since multiple sites were evaluated.

Destination: Candor Chasma/Region G

Coordinates (use UTM NAD27 CONUS): N4251000, W521000

Participants: Kasey (CMD), Vladimir (GHO), and Ben (JOU)

Road(s) and routes per MDRS Map: From the Hab, driving north on Cow Dung Rd to Galileo Rd and Galileo Rd to near area F; walking from area F east to Candor Chasma

Mode of travel: Driving and walking

Operations – December 30th

Crew 236 Operations Report 30-DEC-2021

SOL: 11

Name of person filing report: Pavithra Ravi

Non-nominal systems: 1 radio and 2 suits didn’t function well on EVA.
It is possible they didn’t charge properly as they may have been
unplugged for part of yesterday (during power conservation).

ROVERS

Spirit rover used: Yes
Hours: 148.8
Beginning charge: 100%
Ending charge: 96%
Currently charging: Left unplugged after last EVA to save power. To be
charged by Shannon.

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: Left unplugged after last EVA to save power. To be
charged by Shannon.

Curiosity rover used: Yes
Hours: 157.8
Beginning charge: 100%
Ending charge: 99%
Currently charging: Left unplugged after EVA to save power. To be
charged by Shannon.

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: To be charged by Shannon.

General notes and comments: None

Summary of Hab operations: we received some extra water today (woo!)

WATER USE:
Water (static tank): approx. 130 gallons
Water (loft tank): 40 gallons
Water Meter: 01542971 units
Static to Loft Pump used – yes or no: Yes. It was filled from 12 to 50
gallons this morning.
Static tank pipe heater (on or off): On
Static tank heater (on or off) Off
Toilet tank emptied (no or yes): No

Summary of internet: Nominal

Summary of suits and radios: 1 radio and 2 suits died during the EVA.
It is possible they didn’t charge properly as they may have been
unplugged for part of yesterday (during power conservation).

Summary of GreenHab operations: Nominal
WATER USE: 7 gallons were used today. This leaves about 151 gallons
in the tank.
Heater (on or off): On
Supplemental light (on or off): On
Harvest: 79g cucumber, 14g tomato, 3g basil
Summary of ScienceDome operations: Nominal
Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done
Nothing used.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support:
None

Operations Report- December 20th

Hello! We had a great first day. Thank you for your support!

SOL: 1

Name of person filing report: Pavithra Ravi

Non-nominal systems: Toilet at HAB

Notes on non-nominal systems: Toilet has a leak. A plastic container
is being used to collect the water, which is then sent down the
toilet.

Generator:

Hours run: Information provided by Outpost

From what time last night: Information provided by Outpost

To what time this morning: 7:20 AM

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 55 gallons

Water Meter: 1539024 units

Water (static tank): approx. 415 gallons

Static to Loft Pump used – yes. Refilled at 7:21 PM from 21 to 55 Gallons.

Water in GreenHab: approx. 240.5 gallons

Water in ScienceDome: N/A

Toilet tank emptied: yes, emptied at 11:45 AM

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: Yes

Hours (before EVA 1): 146.3

Beginning charge (Before EVA 1): 100%

Ending charge (On return from EVA 1, before recharging): 99%

Rover was charged before EVA 2.

Hours (before EVA 2): 146.4

Beginning charge (Before EVA 2): 100%

Ending charge (On return from EVA 2, before recharging): 100%

Currently charging: Yes

Opportunity rover used: Yes, for EVA 2

Hours: 76.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, for EVA 1

Hours: 154.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: None

Summary of internet: MOS-MOCC wifi network is no longer available.
Only the regular Mars Hab wifi is in use.

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal. Small concern: the window above
the couch in the Hab does not fit and is letting air in.

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Problems with the dome movement
have been reported to the astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None

Operations Report – December 21st

Crew 236 Operations Report 21Dec2021

SOL: 2

Name of person filing report: Pavithra Ravi

Non-nominal systems:

Notes on non-nominal systems:

Generator:

Hours run: 13 hours 48 minutes

From what time last night: 18:07

To what time this morning: 07:55

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 29 gallons

Water Meter: 1539449 units

Water (static tank): approx. 415 gallons

Static to Loft Pump used: No

Water in GreenHab: approx. 232 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments:

Today the toilet leak worsened, requiring frequent dumping out of the
collected water. This has now been dealt with by keeping the water
valve closed, except for when flushing is required.

The exposed power outlets above the stove in the hab have been covered
with gorilla tape as a safety precaution.

The gaps surrounding the window above the couch in the hab have also
been blocked out with gorilla tape to minimize influx of (lethal)
martian atmosphere into our living area.

The temperature on the upper level of the hab has thankfully been
stabilized at around 70 F. The random bursts of cool air which we
received yesterday seem to have stopped (the trick was keeping the
thermostat at 70 instead of 75).

Summary of internet: Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal.

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the
astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None